OpenStudy (anonymous):

DIff Eq. initial value problem. dy/dt = -y + 5 y(o) = yo

OpenStudy (kc_kennylau):

So dt/dy = 1/(-y+5)

OpenStudy (anonymous):

Why flip it like that?

OpenStudy (kc_kennylau):

So you can now integrate lol

OpenStudy (anonymous):

I dont see hw that helps..sorry I dont understand.

OpenStudy (kc_kennylau):

Integrate 1/(-y+5) lol

OpenStudy (kc_kennylau):

dy of course

OpenStudy (anonymous):

so - ln(-y + 5)

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

So t=-ln(-y+5)

OpenStudy (anonymous):

Is that because we were actually evaluating dy/(-y + 5) = dt. then integrated each side so t = -ln(-y+5)

OpenStudy (kc_kennylau):

Yep you can say that

OpenStudy (anonymous):

okay so t = -ln(-y+5)....

OpenStudy (kc_kennylau):

+C

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

So I need to get this in terms of y?

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

But you find the constant first I think?

OpenStudy (anonymous):

how does e^(-ln(-y+5) simplify?

OpenStudy (kc_kennylau):

\[\Large e^{-\ln(-y+5)}=\frac1{-y+5}\]

OpenStudy (kc_kennylau):

okay I shouldn't have skipped so many steps

OpenStudy (kc_kennylau):

\[\Large e^{-\ln(-y+5)}=\left[e^{\ln(-y+5)}\right]^{-1}=(-y+5)^{-1}=\frac1{-y+5}\]

OpenStudy (anonymous):

ahhh I always wondered how that worked out so nicely like that.

OpenStudy (anonymous):

c wont be part of e's power?

OpenStudy (kc_kennylau):

It would be better to find the value of C before simplifying all that

OpenStudy (anonymous):

How would I do that?

OpenStudy (kc_kennylau):

That's where the initial value is useful

OpenStudy (kc_kennylau):

Substitute the initial value

OpenStudy (anonymous):

for c?

OpenStudy (kc_kennylau):

Well I actually don't know what's y(o)=yo

OpenStudy (anonymous):

Y(0) = Y naught or however it's spelled (;

OpenStudy (kc_kennylau):

What's Y0?

OpenStudy (anonymous):

Means you plug zero into the function y

OpenStudy (kc_kennylau):

but what's y0?

OpenStudy (kc_kennylau):

What you just said was y(0)

OpenStudy (anonymous):

it is y naught aka the inital y value.

OpenStudy (kc_kennylau):

Uh...

OpenStudy (kc_kennylau):

What's the initial value?

OpenStudy (kc_kennylau):

Okay I get it

OpenStudy (kc_kennylau):

So plug that in now

OpenStudy (anonymous):

as in instead of having a integer as the initial value we have a variable name for that integer.

OpenStudy (kc_kennylau):

t=0 and y=y0